OpenStudy (anonymous):

Solve. 4(a + 2) = 14 – 2(3 – 2a) A. all real numbers B. –1 C. no solution D. –2

OpenStudy (solomonzelman):

4(a + 2) = 14 – 2(3 – 2a) expand '4(a+2) and '-2(3-2a)'

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Tell me what you get.

OpenStudy (anonymous):

b

OpenStudy (solomonzelman):

no, that's ain't right.

OpenStudy (anonymous):

c

OpenStudy (solomonzelman):

try to solve for a and tell me if you have any difficulties doing that.

OpenStudy (solomonzelman):

don't guess, use logic.....

OpenStudy (anonymous):

a

OpenStudy (solomonzelman):

\[4(a + 2) = 14 – 2(3 – 2a)\]\[4a + 8 = 14 – 6 + 4a)\]\[4a+8=4a+8---> 4a=4a\]

OpenStudy (solomonzelman):

yes, it is a.

OpenStudy (solomonzelman):

because no matter what you plug in for a, it will work, right?

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

don't get it?

OpenStudy (solomonzelman):

it is a, and I just explained why.

OpenStudy (anonymous):

ohhh ok i get it

OpenStudy (solomonzelman):

4a=4a, or a=a, whatever you plug in for a the equation will be true.

OpenStudy (solomonzelman):

good!

OpenStudy (solomonzelman):

(•‿•)

OpenStudy (anonymous):

thx

OpenStudy (solomonzelman):

Anytime!